OpenStudy (anonymous):

Find the slope if it exists. a.) 2x-4y=8 b.) y=9

OpenStudy (hba):

\[\huge\ Slope=m=\frac{ -coefficient \ of \ x }{ coefficient \ of \ y }\]

OpenStudy (anonymous):

so far i have this: 2x-4y=8 -4y=-2x+8 y=-1/4(-2x+8) y= i get stuck right there

OpenStudy (anonymous):

so far i have this: 2x-4y=8 -4y=-2x+8 y=-1/4(-2x+8) y= i get stuck right there

OpenStudy (phi):

y=-1/4(-2x+8) distribute the -1/4 that means multiply -1/4 times each term inside the parens \[ y= -\frac{1}{4} \cdot -2x -\frac{1}{4} \cdot 8 \] can you simplify this to get it in the form y= mx+b ?

OpenStudy (anonymous):

would you multiply the -1/4*-1/4*8?

OpenStudy (precal):

 for part b) y=9 you can look at the graph to determine the slope

OpenStudy (phi):

would you multiply the -1/4*-1/4*8? no. You have 2 terms. the first term is -1/4 * -2 *x Here is an example with easier numbers 2*3*x is 6*x or 6x you multiply the numbers, but leave the x. now try \[ -\frac{1}{4} \cdot -2 \cdot x\]

OpenStudy (phi):

notice you have a minus times a minus so you get a positive to multiply fractions, multiply top times top and bottom times bottom (whole numbers like 2 have a "bottom" of 1 .... 2 is the same as 2/1 )

OpenStudy (anonymous):

so it would be -1/4 * -1/4 * 8 = 1/16 * 8 = .5 or 1/2

OpenStudy (anonymous):

so my answer would be 1/2???

OpenStudy (phi):

you start with \[ y= -\frac{1}{4} \cdot -2x -\frac{1}{4} \cdot 8 \] when you are done, the x does not disappear. try the first term again: \[ -\frac{1}{4} \cdot -2 \cdot x \]

OpenStudy (anonymous):

sorry this is just all confusing to me...

OpenStudy (anonymous):

.5

OpenStudy (anonymous):

.5x

OpenStudy (phi):

you have just 2 numbers and x multiplied together \[ -\frac{1}{4} \cdot -2 \cdot x\] you can multiply -1/4 times -2 BUT LEAVE the X in

OpenStudy (anonymous):

so it would be .5x?

OpenStudy (phi):

is would be \[ \frac{2}{4} x \] which you can simplify to 1/2 (or 0.5) now to finish this (but you already know the slope, it's the number in front of the x) \[ y= -\frac{1}{4} \cdot -2x -\frac{1}{4} \cdot 8 \] \[ y= \frac{1}{2}x -\frac{1}{4} \cdot 8 \] can you simplify the 2nd term: -1/4 * 8 ?

OpenStudy (anonymous):

-2

OpenStudy (phi):

yes , so the equation is y = 1/2 x - 2 that is in "slope-intercept" form y= mx+b m is the slope (m is the name they give to the number in front of the x) in this problem m is 1/2 the intercept (called b) is -2

OpenStudy (phi):

for question (b) y=9 this is a bit tricky, but we could write this as y = 0x +9 now match this to y= mx+b what is m for this problem?

OpenStudy (anonymous):

0

OpenStudy (phi):

so that is the slope for the line y=9

OpenStudy (anonymous):

ok...thanks!

OpenStudy (anonymous):

is there anyway you could help me with a few more?

OpenStudy (phi):

post the problem

OpenStudy (anonymous):

ok let me close this one